# Disentanglement in a GAN for Unconditional Speech Synthesis

### v1

@@ Abstract
publicationid: pubid: © 2023 IEEE
@@ I Introduction
Direct speech synthesis from a latent space is a very challenging problem, and we therefore restrict ourselves to a limited vocabulary, as was also done in previous work [1, 5, 6]. In this problem setting, recent studies on diffusion models [7] for images [8, 9, 10] has led to major improvements in unconditional speech synthesis. The current best-performing approaches are all based on diffusion modelling [6, 5], which iteratively de-noise a sampled signal into a waveform through a Markov chain [7]. Before this, many studies used generative adversarial networks (GANs) [11] that map a latent vector to a sequence of speech features with a single forward pass through the model. However, performance was limited [1, 2], leading to GANs falling out of favour for this task.
@@
In unconditional speech synthesis experiments, we measure the quality and diversity of the generated samples using objective metrics. We show that ASGAN sets a new state-of-the-art in unconditional speech synthesis on the Google Speech Commands digits dataset [15]. Objective metrics that measure latent space disentanglement indicate that ASGAN has smoother latent representations compared to existing diffusion models. It not only outperforms the best existing models but is also faster to train and faster in inference. Subjective mean opinion scores (MOS) indicate that ASGAN’s generated utterances sound more natural (MOS: 3.68) than the existing best model (SaShiMi [6], MOS: 3.33). We also perform ablation experiments showing intrinsically that our proposed anti-aliasing and adaptive discriminator augmentation techniques are necessary for high-quality and diverse synthesis.
This work is an extension of the conference paper [16], where (apart from the ablation experiments) many of the above intrinsic evaluations were already presented. Here, for the first time, we also present a thorough evaluation of ASGAN’s abilities – we show that ASGAN’s disentangled latent space allows us to perform several tasks unseen during training through simple linear operations in its latent space. Concretely, we demonstrate compelling zero-shot performance on voice conversion, speech enhancement, speaker verification and keyword classification on the Google Speech Commands digits dataset. While not matching the performance of state-of-the-art task-specific systems on all these tasks, our experiments show that a single model designed for disentanglement can achieve reasonable performance across a range of tasks that it hasn’t seen in training. Our work shows the continued competitive nature of GANs compared to diffusion models, and the generalization benefits of designing for disentanglement.
@@ II Related Work
Within unconditional speech synthesis, a substantial body of work focuses on either autoregressive models [23] – generating a current sample based on previous outputs – or diffusion models [5]. Diffusion models iteratively denoise a sampled signal into a waveform through a Markov chain with a constant number of steps [7]. At each inference step, the original noise signal is slightly denoised until, in the last step, it resembles coherent speech. Autoregressive and diffusion models are relatively slow because they require repeated forward passes through the model during inference.
Earlier studies [1, 2] attempted to use GANs [11] for unconditional speech synthesis, which has the advantage of requiring only a single pass through the model. While results showed some initial promise, performance was poor in terms of speech quality and diversity, with the more recent diffusion models performing much better [6]. However, there have been substantial improvements in GAN-based modelling for image synthesis in the intervening years [14, 12, 13]. Our goal is to improve the performance of the earlier GAN-based unconditional speech synthesis models by adapting lessons from these recent image synthesis studies.
Some of these innovations in GANs are modality-agnostic: $R_{1}$ regularization [24] and exponential moving averaging of generator weights [25] can be directly transferred from the vision domain to speech. Other techniques, such as the carefully designed anti-aliasing filters between layers in StyleGAN3 [3] require specific adaptation; in contrast to images, there is little meaningful information in speech below 300 Hz, necessitating a redesign of the anti-aliasing filters.
@@ III-A1 Mapping network
The mapping network $W$ is a simple multi-layer perceptron consisting of several linear layers with leaky ReLU activations between them. As input, it takes in a normally distributed vector $\mathbf{z}\sim Z=\mathcal{N}(\mathbf{0},\mathbf{I})$; in all our experiments we use a 512-dimensional multivariate normal vector, $\mathbf{z}\in\mathbb{R}^{512}$. Passing this vector through the mapping network produces a latent vector $\mathbf{w}=W(\mathbf{z})$ of the same dimensionality as $\mathbf{z}$. As explained in [12], the primary purpose of $W$ is learn to map noise to a linearly disentangled $W$ space, as ultimately this will allow for more controllable and understandable synthesis. $W$ is coaxed into learning such a disentangled representation because it can only linearly modulate channels of the cosine features in each layer of the convolutional encoder (see details below). So, if $\mathbf{w}$ is to linearly shape the speech features, $W$ must learn a mapping that organizes the random normal space $\mathbf{z}$ into one which linearly disentangles common factors of speech variation.
@@ III-A2 Convolutional encoder
The sequence produced by the Fourier feature layer is iteratively passed through Style Blocks, which is based on the encoder layers of the StyleGAN family of models. In each layer, the input sequence and style vector $\mathbf{w}$ are passed through a modulated convolution layer [13]: the final convolution kernel is computed by multiplying the layer’s learnt kernel with the style vector derived from $\mathbf{w}$, broadcasted over the length of the kernel. In this way, the latent vector $\mathbf{w}$ linearly modulates the kernel in each convolution.
@@ IV-C Baseline systems
The autoregressive SaShiMi is originally evaluated using a form of rejection sampling [6] to retain only high probability generated samples for evaluation. We could also perform sampling tricks on all the models (since both diffusion and GAN models also have tractable likelihood measures). However, in the interest of an fair and simple comparison of the inherent performance of each model, we opt to keep the same and most general sampling method for each model (including ASGAN). So, in all experiments, we perform direct sampling from the latent space for the GAN and diffusion models according to the original papers. For the autoregressive models, we directly sample from the predicted distribution for each time-step.
@@ V-A Comparison to baselines
The latent space disentanglement metrics and generation speed are given in Table II. These results are more mixed, with WaveGAN being the fastest model and the one with the shortest latent path length in the $Z$-space. However, this is somewhat misleading since WaveGAN’s samples have low quality (naturalness, as measured by eMOS/MOS) and poor diversity compared to the other models (Table I). This means that WaveGAN’s latent space is a poor representation of the true distribution of speech in the SC09 dataset, allowing it to have a very small path length as most paths do not span a diverse set of speech variation.
@@ VI Solving Unseen Tasks Through Linear Latent Operations
As a reminder, the key aspect of ASGAN’s design is the latent space associated with the vector $\mathbf{w}$ is linearly disentangled. The idea is that, since the $\mathbf{w}$ vector can only linearly affect the output of the model through the Fourier feature layer and modulated convolutions (Fig. 1), $W$ must learn to linearly disentangle common factors of speech variation. If this turns out to be true – i.e. the space is indeed disentangled – then these factors should correspond to linear directions in the $W$ latent space. As originally motivated in studies on image synthesis [12, 13], this would mean that linear operations in the latent space should correspond to meaningful edits in the generated output. In our case, this would mean that if we know the relationship between two utterances, then the linear distance between the latent vectors $\mathbf{w}$ of those utterances should reflect that relationship. E.g. consider a collection of utterances differing only in noise level but otherwise having the same properties. If the space disentangles noise levels, then we should expect the latent points of these utterances to lie in the same linear subspace, reflective of the level of noise.
@@ VI-B1 Style mixing
We can perform voice conversion or speech editing by using style mixing [12] of the latent vectors $\mathbf{w}$. Concretely, we can project two utterances $X_{1}$ and $X_{2}$ to their latent representations $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$. Then – recalling the architecture in Fig. 1 – we can use different $\mathbf{w}$ vectors as input into each Style Block. According to our design motivation of the anti-aliasing filters in Sec. III-D5, the coarse styles are captured in the earlier layers, and the fine styles are introduced in later layers. So we can perform voice conversion from $X_{1}$’s speaker to $X_{2}$’s speaker by conditioning later layers with $\mathbf{w}_{2}$ while still using $\mathbf{w}_{1}$ in earlier layers. This causes the generated utterance to inherent the speaking style (fine) from the target utterance $X_{2}$, but retain the word identity (coarse) from $X_{1}$. By doing the opposite we can also do speech editing: having the speaker from $X_{1}$ say the word in $X_{2}$ by conditioning earlier layers with $\mathbf{w}_{2}$ while keeping $\mathbf{w}_{1}$ in later layers. Furthermore, because the $W$ latent space is continuous, we can interpolate between retaining and replacing the coarse and fine styles to achieve varying degrees of voice conversion or speech editing. We do these style mixing experiments in Sec. VIII-A.
@@ VI-B3 Speaker verification and keyword classification
The previous downstream tasks were generative in nature. The $W$ latent space also allows us to perform discriminative tasks such as speaker verification and keyword classification: determining which speaker or word is present in a given utterance, respectively. For both tasks we use the linear nature of disentanglement: given enrollment utterances containing labelled speakers (for speaker verification) or words (for keyword classification), we invert them to their $\mathbf{w}$ vectors. Then we find the linear projection within the $W$ latent space that maximizes the separation of the labeled characteristic using linear discriminant analysis (LDA). For inference on new data, we invert the input, project it along the LDA axes, and make a decision based on linear distances to other points in the LDA-projected latent space.

### v2

@@ Abstract
publicationid: pubid: © 2024 IEEE
@@ I Introduction
Direct speech synthesis from a latent space is a very challenging problem, with prior efforts only able to model speech in a restricted setting [1, 5, 6]. We therefore also restrict ourselves to the small-vocabulary case here.
In this problem setting, recent studies on diffusion models [7] for images [8, 9, 10] has led to major improvements in unconditional speech synthesis. The current best-performing approaches are all based on diffusion modelling [6, 5], which iteratively de-noise a sampled signal into a waveform through a Markov chain [7]. Before this, many studies used generative adversarial networks (GANs) [11] that map a latent vector to a sequence of speech features with a single forward pass through the model. However, performance was limited [1, 2], leading to GANs falling out of favour for this task.
@@
In unconditional speech synthesis experiments, we measure the quality and diversity of the generated samples using objective metrics. We show that ASGAN sets a new state-of-the-art in unconditional speech synthesis on the Google Speech Commands digits dataset [15]. Objective metrics that measure latent space disentanglement indicate that ASGAN has a more disentangled latent representation compared to existing diffusion models. It not only outperforms the best existing models but is also faster to train and faster in inference. Subjective mean opinion scores (MOS) indicate that ASGAN’s generated utterances sound more natural (MOS: 3.68) than the existing best model (SaShiMi [6], MOS: 3.33). We also perform ablation experiments showing intrinsically that our proposed anti-aliasing and adaptive discriminator augmentation techniques are necessary for high-quality and diverse synthesis.
This work is an extension of the conference paper [16], where (apart from the ablation experiments) many of the above intrinsic evaluations were already presented. Here, for the first time, we profile the generalization benefits of a disentangled latent space. Through an evaluation of ASGAN’s abilities we show that its disentangled latent space allows us to perform several tasks unseen during training through simple linear operations in its latent space. Concretely, we demonstrate compelling zero-shot performance on voice conversion, speech enhancement, speaker verification and keyword classification on the Google Speech Commands digits dataset. While not matching the performance of state-of-the-art task-specific systems on all these tasks, our experiments show that a single model designed for disentanglement can achieve reasonable performance across a range of tasks that it hasn’t seen in training. Our work shows the continued competitive nature of GANs compared to diffusion models, and the generalization benefits of designing for disentanglement.
@@ II Related Work
Within unconditional speech synthesis, a substantial body of work focuses on either autoregressive models [23] – generating a current sample based on previous outputs – or diffusion models [5]. Diffusion models iteratively denoise a sampled signal into a waveform through a Markov chain according to a specified inference schedule [7]. At each inference step, the original noise signal is slightly denoised until, in the last step, it resembles coherent speech. Autoregressive and diffusion models are relatively slow because they require repeated forward passes through the model during inference.
Earlier studies [1, 2] attempted to use GANs [11] for unconditional speech synthesis, which has the advantage of requiring only a single pass through the model. While results showed some initial promise, performance was poor in terms of speech quality and diversity, with the more recent diffusion models performing much better [6]. However, there have been substantial improvements in GAN-based modelling for image synthesis in the intervening years [14, 12, 13].
Palkama et. al. [24] made initial inroads in applying techniques from the earlier StyleGAN models for unconditional speech synthesis. Their main focus, however, was to improve conditional speech synthesis using the digit label to guide generation, with the ultimate goal of building a GAN-based text-to-speech model. Our focus instead is on unconditional speech synthesis, particularly in how a GAN-based approach leads to a disentangled latent space.
@@ III-A1 Mapping network
The mapping network $W$ is a simple multi-layer perceptron consisting of several linear layers with leaky ReLU activations between them. As input, it takes in a normally distributed vector $\mathbf{z}\sim Z=\mathcal{N}(\mathbf{0},\mathbf{I})$; in all our experiments we use a 512-dimensional multivariate normal vector, $\mathbf{z}\in\mathbb{R}^{512}$. Passing this vector through the mapping network produces a latent vector $\mathbf{w}=W(\mathbf{z})$ of the same dimensionality as $\mathbf{z}$. As explained in [12], the primary purpose of $W$ is learn to map noise to a linearly disentangled $W$ space, as ultimately this will allow for more controllable and understandable synthesis. $W$ is coaxed into learning such a disentangled representation because it can only linearly modulate channels of the cosine features in each layer of the convolutional encoder (see details below). So, if $\mathbf{w}$ is to linearly shape the speech features, $W$ should ideally learn a mapping that organizes the random normal space $\mathbf{z}$ into one which linearly disentangles common factors of speech variation.
@@ III-A2 Convolutional encoder
The sequence produced by the Fourier feature layer is iteratively passed through 14 Style Blocks, which are based on the encoder layers of the StyleGAN family of models. In each layer, the input sequence and style vector $\mathbf{w}$ are passed through a modulated convolution layer [13]: the final convolution kernel is computed by multiplying the layer’s learnt kernel with the style vector derived from $\mathbf{w}$, broadcasted over the length of the kernel. In this way, the latent vector $\mathbf{w}$ linearly modulates the kernel in each convolution.
@@ IV-C Baseline systems
ASGAN is a two-stage model (it generates a feature representation, then vocodes the result), while WaveGAN, SaShiMi, and DiffWave are one-stage models (they directly generate the output waveform). Despite SaShiMi+DiffWave being the current top-performing model, our two-stage approach may have an advantage over one-stage approaches because it separates the vocoding task to a separate model.
Lastly, the autoregressive SaShiMi is originally evaluated using a form of rejection sampling [6] to retain only high probability generated samples for evaluation. But, when sampling from a GAN using $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, we also have the exact density value for this sampled $\mathbf{z}$. So, we could perform various sampling tricks on all the models (since both diffusion and GAN models also have tractable likelihood measures). However, in the interest of an fair and simple comparison of the inherent performance of each model, we opt to keep the same and most general sampling method for each model (including ASGAN). So, in all experiments, we perform direct sampling from the latent space for the GAN and diffusion models according to the original papers. For the autoregressive models, we directly sample from the predicted distribution for each time-step.
@@ V-A Comparison to baselines
The latent space disentanglement metrics and generation speed of each model are compared in Table II. These results are more mixed, with WaveGAN being the fastest model and the one with the shortest latent path length in the $Z$-space. However, this is somewhat misleading since WaveGAN’s samples have low quality (naturalness, as measured by eMOS/MOS) and poor diversity compared to the other models (Table I). This means that WaveGAN’s latent space is a poor representation of the true distribution of speech in the SC09 dataset, allowing it to have a very small path length as most paths do not span a diverse set of speech variation.
@@ VI Solving Unseen Tasks Through Linear Latent Operations
As a reminder, the key aspect of ASGAN’s design is that the latent space associated with the vector $\mathbf{w}$ is linearly disentangled. The idea is that, since the $\mathbf{w}$ vector can only linearly affect the output of the model through the Fourier feature layer and modulated convolutions (Fig. 1), $W$ should ideally learn to linearly disentangle common factors of speech variation. If this turns out to be true – i.e. the space is indeed disentangled – then these factors should correspond to linear directions in the $W$ latent space. As originally motivated in studies on image synthesis [12, 13], this would mean that linear operations in the latent space should correspond to meaningful edits in the generated output. In our case, this would mean that if we know the relationship between two utterances, then the linear distance between the latent vectors $\mathbf{w}$ of those utterances should reflect that relationship. E.g. consider a collection of utterances differing only in noise level but otherwise having the same properties. If the space disentangles noise levels, then we should expect the latent points of these utterances to lie in the same linear subspace, reflective of the level of noise.
@@ VI-B1 Style mixing
We can perform voice conversion or speech editing by using style mixing of the latent vectors $\mathbf{w}$. Style mixing is a technique proposed in [12], where they find that StyleGAN models encode coarse information (e.g. facial type, frame positioning) using the $\mathbf{w}$ vector passed to earlier layers, and finer details (e.g. hair color, lighting) using the $\mathbf{w}$ vector passed to later layers. We will qualitatively show that this also applies to speech in ASGAN, where earlier layers control coarse aspects of speech (e.g. digit class) while later layers control finer aspects of speech (e.g. speaker identity).
Concretely, we can project two utterances $X_{1}$ and $X_{2}$ to their latent representations $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$. Then – recalling the architecture in Fig. 1 – we can use different $\mathbf{w}$ vectors as input into each Style Block. According to our design motivation of the anti-aliasing filters in Sec. III-D5, the coarse styles are captured in the earlier layers, and the fine styles are introduced in later layers. So we can perform voice conversion from $X_{1}$’s speaker to $X_{2}$’s speaker by conditioning later layers with $\mathbf{w}_{2}$ while still using $\mathbf{w}_{1}$ in earlier layers. This causes the generated utterance to inherent the speaking style (fine) from the target utterance $X_{2}$, but retain the word identity (coarse) from $X_{1}$. By doing the opposite we can also do speech editing: having the speaker from $X_{1}$ say the word in $X_{2}$ by conditioning earlier layers with $\mathbf{w}_{2}$ while keeping $\mathbf{w}_{1}$ in later layers. Furthermore, because the $W$ latent space is continuous, we can interpolate between retaining and replacing the coarse and fine styles to achieve varying degrees of voice conversion or speech editing. We do these style mixing experiments in Sec. VIII-A. In preliminary experiments we found it optimal to use the first 11 Style Blocks for coarse styles and the remaining 5 Style Blocks for fine styles.
@@ VI-B3 Speaker verification and keyword classification
The previous downstream tasks were generative in nature. The $W$ latent space also allows us to perform discriminative tasks such as speaker verification and keyword classification: determining which speaker or word is present in a given utterance, respectively. For both tasks we use the linear nature of disentanglement: given enrollment utterances containing labelled speakers (for speaker verification) or words (for keyword classification), we invert them to their $\mathbf{w}$ vectors. Then we find the linear projection within the $W$ latent space that maximizes the separation of the labeled characteristic using linear discriminant analysis (LDA). For inference on new data, we invert the input, project it along the LDA axes, and make a decision based on linear distances to other points in the LDA-projected latent space. The idea is to perform a linear projection that maximizes separation based on the chosen characteristic (e.g. speaker identity), otherwise the linear distance could be influenced more by other factors (e.g. noise).